40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3472     Entry time: Wed Aug 25 16:13:32 2010
Author: steve 
Type: Summary 
Category: PEM 
Subject: PSL optical table is back into operation at 32.75" level 

The tile work was done yesterday after noon.

This morning Mike Gerfen and me lowered the enclosure frame to normal height.

Keven- janitor and I removed plastic covers from chambers, racks, SP, MC2 and clean tool boxes.

 

The afternoon Jenne, Kiwamu, Joe and Aiden cleaned the enclosure inside out. The particle count measured zero inside the enclosure with HEPAs on when the covers were

removed. The MOPA and all other components were happy to see us in excellent condition.

This table height is very user friendly!

 

Safety grounds were reconnected.

 

Atm1, new tiles around the concrete slab

Atm2, frame lowered with low cross bars reinstalled

Atm3, the enclosure frame's north west foot is connected to ground

Atm4, PSL optical table is connected to ground at the north east corner through 1 Mohm

Atm5, PSL optical table level at a stimulating, back-friendly height

Attachment 1: P1060764.JPG  1.990 MB  | Hide | Hide all
P1060764.JPG
Attachment 2: P1060768.JPG  2.079 MB  | Hide | Hide all
P1060768.JPG
Attachment 3: P1060777.JPG  2.059 MB  | Hide | Hide all
P1060777.JPG
Attachment 4: P1060776.JPG  2.381 MB  | Hide | Hide all
P1060776.JPG
Attachment 5: P1060780.JPG  2.200 MB  | Hide | Hide all
P1060780.JPG
ELOG V3.1.3-